OpenStudy (nautika27):

1. Larry makes sandwiches for lunch everyday. He can make 6 sandwiches with every 3 tomatoes he buys. If he bought 9 tomatoes, how many sandwiches could he make? 162 9 18 54

OpenStudy (nautika27):

@sammixboo @dan815 @badgilz10 @CountryGurl15

OpenStudy (alexalas):

6/3=?

OpenStudy (nautika27):

btw this lesson is about Ratios and Proportions if that helps !?

OpenStudy (anonymous):

54

OpenStudy (alexalas):

6/3=x/9 cross multiply

OpenStudy (alexalas):

@CODblackops2 don't just give the answer nobody can learn if you just give the answer

OpenStudy (anonymous):

looks like they've got it covered =) keep up the good help yáll

OpenStudy (anonymous):

sure thing

OpenStudy (nautika27):

can u help me with this alexalas